OpenStudy (anonymous):

The plane p has equation 3x+2y+4z=13. A second plane q is perpendicular to p and has equation ax+y+z=4, where a is a constant. (i) Find the value of a. (ii) The line with equation r=j−k+ λ (i+2j+2k)meets the plane p at the point A and the plane q at the point B. Find the length of AB.

OpenStudy (phi):

if 2 planes are perpendicular, their normals are perpendicular the normal of plane 1 is <3,2,4> the normal of plane 2 is <a,1,1>

OpenStudy (anonymous):

and ?

OpenStudy (phi):

do you know how to test if two vectors are perpendicular ? does "dot product" ring any bells?

OpenStudy (anonymous):

yeah

OpenStudy (phi):

can you find a?

OpenStudy (anonymous):

a=2

OpenStudy (anonymous):

huh ?

OpenStudy (phi):

are you saying <3,2,4> dot <2,1,1> = 0 all the numbers are positive.: 6+2+4 does not look promising

OpenStudy (anonymous):

-2

OpenStudy (phi):

the dot product is 0 if the two vectors are perpendicular. with a=2, the dot product is not 0

OpenStudy (phi):

I would do <3,2,4> dot <a,1,1>= 0 3a + 2+ 4 =0 3a+6= 0 3a= -6 a= -2

OpenStudy (anonymous):

yeah ive got it

OpenStudy (phi):

to find the intersection of a line with a plane, start with P=j−k+ a (i+2j+2k) and <3,2,4> dot P = 13 I would write the equation of the line as P= <0,1,-1> + a<1,2,2> and use that in the equation for the plane <3,2,4> dot ( <0,1,-1> + a<1,2,2> ) = 13 solve for a (i.e. lambda, but easier to type) and then find P using the line equation